Calculate the formula weight of a nonelectrolyte, 5 g of which dissolved in 200 g of water gives solution which has a freezing point of negative 1.04 degrees celsius.


1
Expert's answer
2021-03-08T06:13:42-0500

Solution:

1) Determine temperature change (∆T):

The freezing point of pure water is 0°C

∆T = 0°C − (-1.04)°C = 1.04°C


2) Determine how many moles of a nonelectrolyte dissolved:

ΔT = i × Kf × m

Kf for water is 1.86°C kg mol¯1

i = 1 (for nonelectrolyte)

m = the molality of the solute = Moles of nonelectrolyte / kilograms solvent

kilograms solvent = 0.2 kg

Hence,

ΔT = (1) × (1.86°C kg mol¯1) × (Moles of nonelectrolyte / 0.2 kg) = 1.04°C

Moles of nonelectrolyte / 0.2 = 0.55914

Moles of nonelectrolyte = 0.11183 mol


Moles of nonelectrolyte = Mass of nonelectrolyte / Formula weight of nonelectrolyte

Formula weight of nonelectrolyte = Mass of nonelectrolyte / Moles of nonelectrolyte

Hence,

Formula weight of nonelectrolyte = (5 g) / (0.11183 mol) = 44.7 g/mol

Formula weight of nonelectrolyte = 44.7 g/mol


Answer: The formula weight of a nonelectrolyte is 44.7 g/mol.